71. An open-end bond mutual fund is holding a three-year, $1 million face value 5 percent annual coupon bond selling at par.Can someone provide a step-by-step answer key to this question? I am looking forward to hearing from you. Thank you for your patience.
Answer #1

Currently bonds are priced at par. It means, coupon rate = Required rate of return = 5%

A 1% decrease in interest rate means ........ Required rate of return " r " = 5% - 1% = 4%

Price of the bond = Coupon * [ 1 - (1+r)-n ] / r + Face value * ( 1 + r)-n

Coupon = 1000,000 * 5% = 50,000. ........... Note: Even if required return decrease, coupon remains fixed at 5%

r = 0.04

n = 3 Years

= 50000 * [ 1 - (1.04)-3 ] / 0.04 + 1000,000 * (1.04)-3

= 138754.55 + 888996.36

= 1027751

Thus, increase in bond price (or) fund assets = 1027751 - 1000,000 = 27,751